Question 32 3 pts Calculate the pH of an aqueous solution which has a [H3O+] = 3 x 10-M. 0 4.5 0 6.5 O 7.5 O 5.5 8.5
Answer #1

сно*] = 3xiр-н pH = -logo Сно*] log.Cз хіо, 10 -9 t 10 = -0 logo log, : - Lo:474 -41 = -48-523) pH = 8. 5

answered by: ANURANJAN SARSAM